OpenStudy (anonymous):

how to solve 2x^2-5x-3=0 by factoring

OpenStudy (tkhunny):

Well, have you considered factoring it?

OpenStudy (jessiegonzales):

2x^2-5x-3=0 (2x+1)(x-3)=0 Can you solve for X?

OpenStudy (anonymous):

x=-1/2 & 3?

OpenStudy (jessiegonzales):

YES AWESOME :)

OpenStudy (anonymous):

thank you (: can you show me how you did the first step, please?

OpenStudy (anonymous):

@jessiegonzales

OpenStudy (jessiegonzales):

Well i used the formula ax^2+bx+c=0

OpenStudy (anonymous):

how do you factor that? like how did you get (2x+1)(x-3)=0?

OpenStudy (anonymous):

@jessiegonzales

OpenStudy (tkhunny):

2x^2-5x-3 If it is going to factor, it MUST look like this (2x __ ___)(x __ ___) 1 and 2 are the ONLY factors of 2. I has to be what I have written. The rest is a little harder. Fortunately, 1 and 3 are the ONLY factors of 3. Thus, it MUST be one of these. (2x __ 3)(x __ 1) (2x __ 1)(x __ 3) It's also a "-3", so the signs have to be different. This adds more possibilities (2x + 3)(x - 1) (2x + 1)(x - 3) (2x - 3)(x + 1) (2x - 1)(x + 3) That is ALL the possibilities. Just find the one that works. (2x + 3)(x - 1) = 2x^2 + x - 3 -- No good. (2x + 1)(x - 3) = 2x^2 -5x - 3 -- Tripped over it! (2x - 3)(x + 1) = 2x^2 (2x - 1)(x + 3) = 2x^2 You can try the last two if you like, just to prove they don't work.